OpenStudy (sugarmochi):

Help ugh Solve for x given (x+4)^4 - 5(x+4)^2+6=0

OpenStudy (3mar):

May I help?

OpenStudy (sugarmochi):

PLEASE DO ! ! !

OpenStudy (3mar):

Thank you.

OpenStudy (3mar):

Hint: put \[y=(x+4)\] and tell me what you got?

OpenStudy (sugarmochi):

(y^4) - 5(y^2) + 6 = 0

OpenStudy (3mar):

Great! \[y^4-5y^2+6=0\] If you can, it'b better to use the equation box to type in your expressions. Now can you factor this equation into two factors firstly?

OpenStudy (3mar):

Need assistance?

OpenStudy (sugarmochi):

would it be \[(y ^{2} - 2) (y ^{2}\pm 3)\]

OpenStudy (3mar):

why \[\pm \]?

OpenStudy (sugarmochi):

because for the previous equation i got \[y ^{2} (y ^{2}-3) -2(y ^{2}+3)\]

OpenStudy (3mar):

Man, it is just like that! \[(y^2-2)(y^2-3)\] and make a check by expand it

OpenStudy (sugarmochi):

ohhhh okay okay

OpenStudy (3mar):

After that you can put your touch by replacing y and plug in its substitution that you have set at the beginning!

OpenStudy (sugarmochi):

( \[(x^4+14) (x^4-13)\]

OpenStudy (sugarmochi):

is it- ?

OpenStudy (3mar):

What I meant is to replace y with (x+4) so if you replace y^2, you will get (x+4)^2 in other words, the final result will be \[[(x+4)^2-2][(x+4)^2-3]\] got it?

OpenStudy (sugarmochi):

ahh yeah

OpenStudy (3mar):

Is it clear now?

OpenStudy (sugarmochi):

im still not sure how to go on from there-

OpenStudy (3mar):

Where are you stuck?

OpenStudy (sugarmochi):

i still dont really know how to find x

OpenStudy (3mar):

Oh sorry I thought that the question just to factor it. Ok. As you got the final result we have, can you equate every parentheses to zero?

OpenStudy (sugarmochi):

Equate parentheses ?

OpenStudy (3mar):

\[(x+4)^2-3=0\] and \[(x+4)^2-2=0\]

OpenStudy (sugarmochi):

ohhhh

OpenStudy (3mar):

I think you got it!

OpenStudy (sugarmochi):

yeah yeah i do! Is it possible for you to help me with another question-

OpenStudy (3mar):

Of course, brother! Why not? Are you satisfied of that one firstly?

OpenStudy (sugarmochi):

Yeah i managed to get the answers thnks ! Can you also help me w the question Solve for x given \[(x^2-2x)^2 - 11(x^2-2x)+24=0\]

OpenStudy (3mar):

Can we start in new post, please? I'd appreciate!

OpenStudy (sugarmochi):

oky doky

OpenStudy (3mar):

Thanks